When incident wavelength is lamda, stopping potential is 3 Vo. If incident wavelength is 2 lambda then stopping potential is Vo. Find out threshold wavelength in terms of lambda

Grade:11

1 Answers

Adarsh
6 years ago
Dear Preeti ,
 
Let the threshold wavelength be λ0 ,
Now A/Q           3V0 = hc/(λ– λ)   ….....  ( 1 )
                        
                           V0= hc/(λ– 2λ)  …....... ( 2 ) 
 
dividing 1 by 2 we get  ,  3 = (λ– 2λ) / (λ– λ)    => λ0 = λ/2 .
 
Hence the threshold wavelength is half of lambda..
 
Thanks and Regards
Adarsh